# Cyber Attacks Detection, Prevention, and Source Localization in Digital Substation Communication using Hybrid Statistical-Deep Learning

### v1

@@ Abstract
The digital transformation of power systems is accelerating the adoption of IEC 61850 standard. However, its communication protocols, including Sampled Values (SV), lack built-in security features such as authentication and encryption, making them vulnerable to malicious packet injection. Such cyber attacks can delay fault clearance or trigger unintended circuit breaker operations. While most existing research focuses on detecting cyber attacks in digital substations, intrusion prevention systems have been disregarded because of the risk of potential communication network disruptions. This paper proposes a novel method using hybrid statistical-deep learning for the detection, prevention, and source localization of IEC 61850 SV injection attacks. The method uses exponentially modified Gaussian distributions to model communication network latency and long short-term memory and Elman recurrent neural network to detect anomalous variations in the estimated probability distributions. It effectively discards malicious SV frames with minimal processing overhead and latency, maintains robustness against communication network latency variation and time-synchronization issues, and guarantees a near-zero false positive rate in non-attack scenarios. Comprehensive validation is conducted on three testbeds involving industrial-grade devices, hardware-in-the-loop simulations, virtualized intelligent electronic devices and merging units, and high-fidelity emulated communication networks. Results demonstrate the method’s suitability for practical deployment in IEC 61850-compliant digital substations.
Index Terms:
Cyber Attacks, Deep Learning, Digital Substations, IEC 61850 Sampled Values, Intrusion Detection and Prevention System, Statistical Analysis.
@@ I Introduction
As proved by the latest cyber attacks on Ukraine’s power grid in 2015, 2016, and 2022 [1, 2, 3], and the attempted one on the United Kingdom’s power grid in 2020 [4], it is crucial to provide Industrial Control Systems (ICS), and, in broader terms, Operational Technology (OT) infrastructures, with state of the art cyber security controls to ensure a secure and resilient power system operation. The growing integration of Information Technology (IT) and OT infrastructures enables efficient power system operation and management. However, it also expands the attack surface and increases the risk of cyber attacks by introducing new cyber security threats [5].
@@
Communication protocols such as those defined in IEC 61850 standard have been developed for advanced protection, automation and control in digital substations. IEC 61850 notably lacks fundamental network security features such as authentication and encryption. This deficiency is especially evident in the IEC 61850 Sampled Values (SV) protocol, where the absence of robust security mechanisms allows malicious actors to inject crafted packets into the substation OT communication network, potentially leading to major disruptions such as delayed fault clearance or unintended circuit breaker operations. Furthermore, these OT disruptions can affect power system stability, cause cascading failures and lead to power outages or even a complete blackout [6], [7].
A viable option for defending against cyber attacks on digital substations is the deployment of network and host-based Intrusion Detection and Prevention Systems (IDS and IPS, respectively). Whereas a plethora of different IDSs for smart grids have been proposed in the literature [8], the deployment of IPSs in digital substations has been disregarded due to the general reluctance to implement such solutions in OT infrastructures. This reluctance arises from the risk of potential OT communication network disruptions, increased communication latencies, packet drops, and required computational resources. In this paper, a novel method is proposed for the detection, prevention, and source localization of IEC 61850 SV-based cyber attacks on digital substations.
The contributions of this paper are summarized as follows:
@@
    We present the first integrated system for intrusion detection, prevention, and attack source localization to protect against IEC 61850 SV injection attacks in digital substations. The method leverages statistical features of SV frames arrival times. The Exponentially Modified Gaussian (EMG) distribution is used to accurately model OT communication network latency variations and devices clock drift during time synchronization loss holdover conditions.
@@
    We propose a real-time hybrid method that combines statistical modeling with Deep Learning (DL) to identify and mitigate all known types of SV injection attacks. The proposed method ensures a near-zero False Positive Rate (FPR) when no cyber attacks are conducted, introduces negligible additional communication latency, and provides a throughput of more than 100,000 SV frames per second. This makes the proposed method suitable for deployment in digital substations. Furthermore, the proposed method is resilient to communication network latency variations and time-synchronization issues among devices.
@@
    We extend the core method to detect and localize the compromised Intelligent Electronic Devices (IEDs) within High-availability Seamless Redundancy (HSR) rings. This is achieved by analyzing the spatial-temporal correlations in the Probability Distribution Functions (PDFs) of the SV frames arrival times across multiple IEDs.
The proposed method is validated and evaluated on three IEC 61850-compliant testbeds including industrial-grade devices, i.e., network switches, IEDs, Merging Units (MU), Hardware-in-the-Loop (HIL) simulations, virtualized IEDs (vIEDs), virtualized MU (vMU), and high-fidelity emulated substation OT communication network and components.
The remainder of the paper is organized as follows. In Section II, background information about the SV protocol and its cyber security concerns are provided. In Section III, state-of-the-art solutions for the detection and prevention of SV injection attacks are discussed and compared with the method proposed in this work. The proposed method is described in Section IV and then extensively validated and evaluated in Section V. Finally, Section VI concludes the paper.
@@ II-A IEC 61850 Sampled Values
IEC 61850 SV is a publisher-subscriber protocol used for three-phase current and voltage measurements reporting from MUs to IEDs. MUs publish SV frames in multicast, whereas IEDs process only the SV streams to which they are subscribed. The protocol was initially introduced in IEC 61850-9-2, further refined in IEC 61850-9-2LE, and ultimately standardized in IEC 61869-9. Various SV frame publishing rates are specified based on system frequency and application requirements. In digital substations, the most used rates are 4000 and 4800 frames per second for 50 Hz and 60 Hz power systems, respectively. Additionally, the standard defines the structure of SV frames, including the smpCnt field, a counter that increments with each published frame and resets to zero every second. Consequently, the smpCnt value ranges from zero to the number of frames published each second minus one. This part of the standard provides a certain level of determinism on the specific arrival time of each of the SV frames marked with a specific smpCnt value. In fact, given a specific second $i$, a predefined number of frames per second $FS$, and a smpCnt field value $c$, the theoretical frame arrival time $F_{a}^{e}(i,c)$ is computed as follows:
@@
As it will be discussed in Section IV, the determinism in the frames’ arrival time gives an information advantage to discern legitimate and malicious SV frames, detect MitM attacks, and locate compromised IEDs within the OT communication network. To further increase the determinism of the SV frames arrival time, as recommend in IEC 61850-90-4, priority tags and proper communication network engineering should be used to ensure the lowest possible latency for SV frames [9].
@@ II-B Sampled Values Cyber Security
Considering the strict latency requirements of power system protection schemes of 3 ms, the IEC 62351 standard stipulates against the application of digital signatures due to increased latency and processing times. To counteract most of the possible cyber attacks exploiting the SV protocol, the standard recommends the usage of Message Authentication Codes (MAC) relying on HMAC or AES-GMAC to provide message authentication. While appending the computed MAC to published frames helps mitigate replay and spoofing attacks, it also introduces multiple challenges and limitations. These include the increased device computational load, communication latency, and message overhead, need for Public Key Infrastructure (PKI) and Key Distribution Center (KDC) deployment, lack of support for legacy devices not supporting IEC 62351-6, and risk of pre-shared keys violation. Moreover, each device belonging to the same Group Domain of Interpretation (GDOI) has access to the same shared secret key, and thus can perform spoofing attacks.
@@
Summarizing, once an attacker is able to gain access to the digital substation OT communication network [5], the cyber attacks that can be launched against an SV subscriber include: (1) flooding, where a large number of SV frames are injected into the communication network by the attacker to exhaust subscriber resources; (2) spoofing, where the attacker injects malicious SV frames pretending of being the legitimate SV publisher; (3) replay, where the attacker re-injects into the network previously sniffed SV frames; (4) high smpCnt attack, in which an attacker injects an SV frame with a high smpCnt filed value, causing the SV subscriber implementing the replay protection mechanism as mandated in IEC 62351-6 to discard any further legitimate SV frame with a lower smpCnt; (5) Man-in-the-Middle (MitM), where an attacker takes control of a device forwarding SV frames from the publisher to the subscriber and tamper with the forwarded SV frames; (6) access to the pre-shared secret key, in which the attacker gains access to the pre-shared secret key used for MAC generation and validation.
@@
In this paper, to evaluate the effectiveness of the proposed method, the performed cyber attacks assume the injection of malicious SV frames which are completely indistinguishable from the legitimate ones. In other words, the Hamming distance between the content of legitimate and malicious SV frames is equal to zero. By proving the method effectiveness in this worst-case scenario, the results can be extended to cases where false measurements are injected, thus leading to a Hamming distance greater than 0.
III Related Work
To address the security concerns presented in the previous sections, several works proposed and evaluated the adoption of authentication and/or encryption for GOOSE and SV [10, 11, 12, 13, 14]. However, the usage of cryptographic schemes raises challenges and limitations as discussed in Section II. Other works proposed novel methodologies for intrusion detection considering the measurements reported in the SV frames. In [15], several Machine Learning (ML) algorithms are tested to detect False Data Injection (FDI) attacks by analyzing the measurements contained in the SV frames and distinguishing between normal and abnormal micro-grid and substation behaviors. Even the most promising solution, which relies on extreme randomized trees, provides good but not impressive performances, with an F1-score of around 91% depending on the different FDI attacks and fault combinations.
In [16], an algorithm based on immune system of negative selection and self-shape optimization is introduced to distinguish between busbar faults and SV attacks. The proposed detector focuses on assessing the validity of the measurements received and provides promising detection performance with a True Positive Rate (TPR) and FPR of around 93% and 7%, respectively, but causes an operation delay ranging from 0.48 to 9.21 ms. Another approach is described in [17], where a neural network-based forecaster is used to detect spoofed SV frames. Whenever a replay attack is detected by checking the smpCnt field value, the frame is sent to the anomaly detector which, using the latest 20 measurements received through the SV stream, estimates the next expected measurement, and if the difference between the forecasted measurement value and the received one is above a certain threshold, the frame is marked as malicious and discarded. A key limitation of this type of IDS is that relying on reported measurements can be misleading during replay attacks where normal operating conditions are injected. In such cases, the measurements appear consistent with the physical model used for validation. To overcome this limitation, other authors proposed IDSs considering network traffic and SV frames information. These solutions rely on information such as the number of SV frames received each second ($FS$), the received frames’ smpCnt, the Inter-Frame Arrival Time (I-FAT), and the frames’ arrival time ($F_{a}$).
In [18], a solution to detect FDI attacks on GOOSE and SV is presented. For what concerns the SV stream injection part, the proposed solution checks that the smpCnt field of the received SV frames is always increasing. The time required to perform the replay attack protection is equal to 0.013 ms, which is below the inter-frame arrival delay of 0.21 ms, and comparable to the processing times of other solutions (0.006, 0.29, 0.049 ms for the solutions proposed in [13], [17], and [15], respectively). However, this type of solution is not enough to detect advanced SV injection attacks, e.g., MitM. In [19], a feed-forward neural network is used to correlate two SV stream parameters, i.e., $FS$ and I-FAT, with the SV publisher that published the SV stream in the network. The proposed method achieves high accuracy in identifying the source publisher of a received stream. However, this performance is attributed to the fact that the second publisher did not attempt to replicate the identity of the first. If impersonation was attempted, a malicious publisher could emulate the SV stream parameters used by the classifier, potentially bypassing detection.
In [20] and [21], a network-based Anomaly Detection System (ADS) to detect malicious activities of GOOSE and SV across substations is presented and then further expanded. For what concerns SV, the system relies on three violation control methods to detect the anomalies. The first violation method checks the number of SV frames received in the last second, and if this value is higher than the threshold value, the anomaly is identified. The second method offers replay protection by checking that the smpCnt field value is always increasing. Finally, the last check verifies that the SV stream identifier and name did not change, implying that the MU configuration did not change. When one of these three methods detects an anomaly, an alert is sent to the operator and a disconnect control command is sent to the firewall to block the intruder’s connection. Even though the performed simulation results confirm excellent detection performance, with FPR and False Negative Rates (FNR) equal to 0.013% and 0.02%, respectively, it is unclear how the firewall can block intruder’s SV injected stream, but not the original one, given that it is not possible to discern the original and injected SV frames. This is of crucial importance, otherwise for the attacker it would be enough to send one single SV frame to the ADS to block the original SV stream reception. Also, if an attacker can block the original SV stream and inject its own, the proposed ADS wouldn’t be able to detect the attack because the number of SV frames received by the IED will not be more than the set threshold, the counter of the injected frames can be easily set by the attacker as always increasing, and the stream identifier can be matched with the one of the original SV stream published by the legitimate MU.
TABLE I: Comparison of the discussed solutions in terms of information used for attack detection. (Y=Yes, N=No, P=Partial).
PHY FS smpCnt I-FAT $\mathbf{F_{a}}$
Ustun et al. [15] Y N N N N
Mo et al. [16] Y N N N N
El Hariri et al. [17] Y N Y N N
Hussain et al. [18] Y N Y N N
Wannous et al. [19] N Y N Y N
Hong et al. [20, 21] N Y Y N N
Eynawi et al. [22] Y N N Y Y
Delhomme et al. [23] N N Y Y Y
Manzoor et al. [24] Y N Y N Y
Hong et al. [25] N Y Y N N
Our Solution N Y Y Y Y
TABLE II: Discussed Solutions’ Protection Capability Against Different Types of Attacks. (Y = Yes, N = No, P = Partial).
(1) (2) (3) (4) (5) (6)
Ustun et al. [15] N Y N N P Y
Mo et al. [16] N Y N N P Y
El Hariri et al. [17] N N Y Y P Y
Hussain et al. [18] P N Y N N Y
Wannous et al. [19] N N N N N Y
Hong et al. [20, 21] P N Y N N Y
Eynawi et al. [22] N N Y N N Y
Delhomme et al. [23] P N N N N Y
Manzoor et al. [24] N N Y Y N Y
Hong et al. [25] P N Y Y N Y
Our Solution P Y Y Y Y Y
In [25], a collaborative intrusion detection and prevention system for GOOSE and SV is presented. For SV, the number of received SV frames and the consistency of the smpCnt parameter are checked. When an anomaly is detected, the protection scheme is disabled to prevent unintended circuit breaker operations and thus leading to an implicit Denial of Service (DoS) caused by the IPS. In [22], a ML-based feature selection algorithm for detecting attacks targeting MMS and SV protocols is proposed. After feature selection, the tested Random Forest (RF)-based IDS provides an accuracy of 99.91% in detecting standard SV injection attacks. Still, the methodology fails in detecting more advanced attacks, e.g., replay with spoofing or MitM. In [23], different ML models for the detection of DoS attacks are presented; still, even the most effective model, that is RF, provides an accuracy of 84.12%. Finally, in [24], an approach leveraging the in-context learning ability of transformer architecture for the detection of novel attacks against IEC 61850 protocols is introduced. The proposed methodology reaches an accuracy close to 100% for the detection of SV injection and replay attacks, but results are not provided for the detection of advanced attacks, e.g., spoofing and MitM. Moreover, the time required for detecting the attacks ranges from 2,44 to 100 ms depending on the available computation power. This does not allow the methodology to be applied for real-time intrusion prevention purposes.
In Tables I and II, the solutions proposed in the literature are compared in terms of information used for intrusion detection, and protection capabilities against the attacks described in Section II. It is worth mentioning that, contrary to our solution, none of the solutions proposed in the literature can identify the source of the SV attack and allow the protection scheme to operate without any additional negligible delays during a cyber attack.
IV Cyber Attacks Detection, Prevention, and Source Localization

### v2

@@ Abstract
The digital transformation of power systems is accelerating the adoption of IEC 61850 standard. However, its communication protocols lack built-in authentication and encryption, leaving them vulnerable to Man-in-the-Middle (MitM) and malicious frame injection attacks that can disrupt protection schemes operation. While most existing research focuses on detecting cyber attacks in digital substations, intrusion prevention systems have been largely overlooked due to concerns about potential network disruptions. To address this gap, this paper proposes an integrated hybrid statistical-deep learning method for detecting, preventing, and localizing IEC 61850 Sampled Values (SV)-based cyber attacks. The method models SV frames arrival times using exponentially modified Gaussian distributions and prevents malicious frames from reaching the targeted Intelligent Electronic Devices (IEDs). Malicious frames are dropped with minimal processing overhead and latency, while the method remains robust to network latency, jitter, and time-synchronization issues, and ensures a near-zero false positive rate under non-attack conditions. Long short-term memory and Elman recurrent neural networks are used to identify anomalous variations in the estimated probability distributions across IEDs for detecting and localizing MitM attacks on SV streams. The proposed method is validated across three testbeds comprising industrial-grade communication and protection devices, hardware-in-the-loop simulations, virtualized IEDs and merging units, and high-fidelity emulated networks. Results demonstrate the method’s practicality and effectiveness for deployment in IEC 61850-compliant digital substations.
@@ I Introduction
As proved by the latest cyber attacks on Ukraine’s power grid in 2015, 2016, and 2022 [3, 25, 18], and the attempted one on the United Kingdom’s power grid in 2020 [5], it is crucial to provide Industrial Control Systems (ICS), and, in broader terms, Operational Technology (OT) infrastructures, with state-of-the-art cyber security controls to ensure a secure and resilient power system operation. The growing integration of Information Technology (IT) and OT infrastructures enables efficient power system operation and management. However, it also expands the attack surface and increases the risk of cyber attacks by introducing new cyber security threats [4].
@@
Communication protocols defined in IEC 61850 standard have been developed to support advanced protection, automation, and control in digital substations. However, IEC 61850 inherently lacks fundamental network security mechanisms such as authentication and encryption. This deficiency is especially evident in the IEC 61850 Sampled Values (SV) protocol, where the absence of cryptographic protection allows malicious actors to inject forged frames into the substation OT communication network, potentially leading to major disruptions such as delayed fault clearance or unintended circuit breaker operations. Furthermore, these OT disruptions can affect power system stability, cause cascading failures and lead to large-scale outages or even a complete blackout [23, 22].
Several works in the literature have proposed the adoption of authentication and encryption to secure SV frames [15, 11, 8, 24, 26]. However, cryptographic schemes raise several challenges and limitations, including additional computational load, communication latency, message overhead, and increased infrastructure complexity. Another viable option to enhance cyber security in digital substation is the deployment of Intrusion Detection and Prevention Systems (IDSs and IPSs, respectively) [10]. Multiple studies have proposed intrusion detection methods based on the current and voltage measurements reported in SV frames [27, 20, 7, 16]. Nonetheless, a key limitation of these IDSs is that relying solely on reported measurements can be misleading during replay attacks, in which normal operating conditions are re-injected by the attacker. In such cases, the measurements appear consistent with the physical model used for validation, and SV attack detection fails. To overcome this limitation, other authors have proposed IDSs that consider communication network traffic and SV frames header, relying on information such as the number of SV frames received each second ($FS$), the value of the smpCnt field, and the Inter-Frame Arrival Time (I-FAT) [14, 6, 16, 28, 12, 9, 19, 13]. While these latest detection methods can effectively identify basic SV injection attacks, they fail against more sophisticated attacks, such as replay with spoofing and Man-in-the-Middle (MitM). In particular, when an attacker compromises a forwarding device, blocks the legitimate SV stream, and injects a malicious one, detection becomes infeasible because the attacker can simultaneously maintain the expected SV frames publishing rate ($FS$), increment the smpCnt field consistently, spoof the SV stream identifier, and closely match the expected inter-frame arrival timing. Moreover, the time required to detect the attacks ranges from a few hundred microseconds [27] to hundreds of milliseconds [19], depending on the available computational power and the implemented detection method. For these reasons, existing approaches cannot be applied for real-time intrusion prevention, MitM attacks detection, and malicious device localization. Finally, none of the existing works exploits the informative advantage provided by the statistical properties of the SV frames arrival time shift ($F_{as}$) to detect, prevent, and localize SV-based injection attacks, which is one of the contributions of this paper, as summarized in Table I.
TABLE I: SV frames features used for cyber attacks detection.
PHY FS smpCnt I-FAT $\mathbf{F_{as}}$
Ustun et al. [27]
Mo et al. [20]
El Hariri et al. [7]
Hussain et al. [16]
Hong et al. [14]
Delhomme et al. [6]
Wannous et al. [28]
Hong et al. [13, 12]
Eynawi et al. [9]
Manzoor et al. [19]
Our Solution
While existing works focused on intrusion detection, the deployment of IPSs in digital substations has so far received limited attention, mainly due to a general reluctance to adopt such solutions in OT infrastructures. This reluctance primarily stems from the risk of potential disruptions to time-critical OT communication networks in digital substations, including increased latency, packet loss, and additional computational overhead. Nevertheless, detection alone is not sufficient to ensure resilient power system operation. Even when an attack is promptly identified, its impact on protection and control functions may already have occurred. Therefore, proactive intrusion prevention mechanisms are essential to mitigate or entirely avoid undesired consequences before they propagate through the OT system, while ensuring no interference with protection, automation, and control functionalities.
To address these gaps, this paper proposes an integrated method for detecting, preventing, and localizing the source of IEC 61850 SV cyber attacks in digital substations, aiming to enhance both situational awareness and operational security while meeting real-time communication requirements for protection schemes. The main contributions of this paper are summarized as follows:
@@
    We present the first integrated method combining statistical modeling and Deep Learning (DL) for real-time detection, prevention, and source localization of IEC 61850 SV cyber attacks in digital substations. The method introduces a novel analytical formulation of SV frames arrival times statistics using Exponentially Modified Gaussian (EMG) probability distributions to accurately model network latency, jitter, and device clock drift during time synchronization loss. The combination of the statistical approach with DL enables the identification and mitigation in real-time of MitM and all known techniques of SV injection attacks.
@@
    We employ Long Short-Term Memory (LSTM) models to capture temporal deviations in the estimated probability distributions, enabling timely detection of ongoing MitM attacks on SV streams against the protected Intelligent Electronic Devices (IEDs).
@@
    We adopt Elman Recurrent Neural Networks (RNNs) to combine and correlate the estimated probability distribution across multiple IEDs within High-availability Seamless Redundancy (HSR) rings. By analyzing the spatio-temporal correlations among Probability Distribution Functions (PDFs), the compromised device used to perform the MitM attack on SV streams can be precisely localized.
  4. 4.
    We utilize the estimated PDFs to discern legitimate and malicious frames in real-time, allowing malicious frames to be discarded before affecting IED protection logic. This provides an effective intrusion prevention mechanism with minimal processing overhead and high operational explainability.
  5. 5.
    We validate the proposed method for real-time detection, prevention, and source localization of IEC 61850 SV cyber attacks in digital substation on two lab testbeds and an industrial testbed at Elia transmission system operator in Belgium comprising industrial-grade protection and communication devices, hardware-in-the-loop simulations, virtualized IEDs (vIEDs) and merging units (vMUs), and high-fidelity emulated OT networks. Results demonstrate accurate detection and localization of MitM attacks on SV streams, reliable filtering of forged SV frames, a near-zero False Positive Rate (FPR) in the absence of cyber attacks, negligible additional latency, and throughput exceeding 100,000 SV frames per second. The method remains resilient to communication latency, jitter, and time synchronization deviations, making it suitable for deployment in IEC 61850-based digital substations.
The remainder of the paper is organized as follows. In Section II, background information about the SV protocol and cyber security concerns is provided. In Section III, the integrated method proposed for the detection, prevention, and source localization of IEC 61850 SV-based cyber attacks is presented. In Section IV, the proposed method is extensively validated, evaluated, and compared with the related state-of-the-art solutions. Finally, Section V concludes the paper.
@@ II Background
As introduced in the previous sections, the operation of digital substations relies on time-critical communication services defined within the IEC 61850 standard. Understanding both the operational principles of the SV protocol and its associated cyber security requirements is essential for assessing vulnerabilities and designing effective defence mechanisms. The following subsections first provide an overview of the IEC 61850 SV protocol and its deterministic timing characteristics, followed by a discussion on the security challenges and countermeasures proposed in the relevant IEC standards.
@@ II-A IEC 61850 Sampled Values
IEC 61850 SV is a publisher-subscriber protocol used for three-phase current and voltage measurements reporting from Merging Units (MUs) to IEDs. MUs publish SV frames in multicast, whereas IEDs process only the SV streams to which they are subscribed. The protocol was initially introduced in IEC 61850-9-2, further refined in IEC 61850-9-2LE, and ultimately standardized in IEC 61869-9. Various SV frame publishing rates are specified based on system frequency and application requirements. In digital substations, the most used rates are 4000 and 4800 frames per second for 50 Hz and 60 Hz power systems, respectively. Additionally, the standard defines the structure of SV frames, including the smpCnt field, a counter that increments with each published frame and resets to zero every second. Consequently, the smpCnt value ranges from zero to the number of frames published each second minus one. This part of the standard provides a certain level of determinism on the specific arrival time of each of the SV frames marked with a specific smpCnt value. In fact, given a specific second $i$, a predefined number of frames per second $FS$, and a smpCnt field value $c$, the theoretical frame arrival time $F_{a}^{e}(i,c)$ is computed as follows:
@@
This theoretical determinism in the frames’ arrival time across publishing periods provides an information advantage to discern legitimate and malicious SV frames, detect MitM attacks, and locate compromised IEDs within the OT communication network. To further increase the determinism of the SV frames arrival time, as recommended in IEC 61850-90-4, priority tags and proper communication network engineering should be used to ensure the lowest possible latency and jitter for SV frames [29].
@@ II-B Sampled Values Cyber Security
Considering the strict latency requirements of power system protection schemes of 3 ms, the IEC 62351 standard stipulates against the application of digital signatures due to increased latency and processing times. To counteract most of the possible cyber attacks exploiting the SV protocol, the standard recommends the usage of Message Authentication Codes (MAC) relying on Hash-based MAC (HMAC) or Advanced Encryption Standard-Galois MAC (AES-GMAC) to provide message authentication. While appending the computed MAC to published frames helps mitigate replay, spoofing, and MitM attacks, it also introduces multiple challenges and limitations. These include the increased publishing and subscribing devices computational load, communication latency, and message overhead, the need for Public Key Infrastructure (PKI) and multiple Key Distribution Centers (KDCs) deployment, lack of support for legacy devices not supporting IEC 62351-6, and the risk of pre-shared keys violation. Moreover, each device belonging to the same Group Domain of Interpretation (GDOI) has access to the same shared secret key, and thus can perform spoofing attacks.
@@
Summarizing, once an attacker is able to gain access to the digital substation OT communication network [4], the cyber attacks that can be performed against an SV subscriber include: (1) flooding, where a large number of SV frames are injected into the communication network by the attacker to exhaust subscriber resources; (2) spoofing, where the attacker injects malicious SV frames pretending of being the legitimate SV publisher; (3) replay, where the attacker re-injects into the communication network the previously sniffed SV frames; (4) high smpCnt attack, in which an attacker injects one or multiple SV frames with a high smpCnt field value, causing the SV subscriber implementing the replay protection mechanism as mandated in IEC 62351-6 to discard any further legitimate SV frame with a lower smpCnt; (5) MitM, where an attacker takes control of a device forwarding SV frames from the publisher to the subscriber and tampers with the forwarded SV frames; (6) access to the pre-shared secret key, in which the attacker gains access to the pre-shared secret key used for MAC generation and validation and injects authenticated malicious SV frames.
@@
In this paper, to evaluate the effectiveness of the proposed method, the performed cyber attacks assume the injection of malicious SV frames that are completely indistinguishable from the legitimate ones. In other words, the Hamming distance between the content of legitimate and malicious SV frames is equal to zero. By proving the method’s effectiveness in this worst-case scenario, the results can be extended to cases where false measurements are injected, thus leading to a Hamming distance greater than 0.
III Cyber Attacks Detection, Prevention, and Source Localization